OpenStudy (anonymous):

which answer is the solution of the equation below, rounded to the nearest tenth 2.1x-4.09=6.98x+1.73 -0.5 -1.2 -1.1 -0.4

OpenStudy (anonymous):

move the variables to one side and the real numbers to the other side I got -5.82=9.08x

OpenStudy (anonymous):

i think......

OpenStudy (anonymous):

oops i messed up should be -5.82=4.88x

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

yes, it is b x would round out to 1.192622

OpenStudy (anonymous):

-4.88x-4.09=1.73 -4.88x=5.82 Divide each term in -4.88x=5.82 by -4.88 x=-1.19262295 Rounded to the nearest tenth x=-1.2

OpenStudy (anonymous):

@genaricstudier